(Cos(2x))^1/(x^2) please help with the derivative. Thanks
1
Expert's answer
2021-08-18T14:20:39-0400
"y=(\cos(2x))^{1/x^2}"

"\ln y=\dfrac{\ln(\cos(2x))}{x^2}"

Differentiate both sides with respect to "x"


"\dfrac{y'}{y}=\dfrac{-2x^2\tan(2x)-2x\ln(\cos(2x))}{x^4}"

"y'=-\dfrac{2}{x^3}\big(x\tan(2x)+\ln(\cos(2x))\big)\big(\cos(2x)\big)^{1/x^2}"